A rectangular garden is to be surrounded by a walkway of constant width. The garden’s dimensions are 30 ft by 40 ft. The total area, garden plus walkway, is to be 1800 ft2. What must be the width of the walkway to the nearest thousandth?
Found 2 solutions by checkley75, renevencer22:
Answer by checkley75(3666)

(30+2X)(40+2X)=1800
1200+80X+60X+4X^2=1800
4X^2+140X+1200-1800=0
4X^2+140X-600=0
X^2+35X-150=0 USING THE QUADRATIC EQUATION WE GET:
X=(-35+-SQRT[35^2-4*1*-150])/2*1
X=(-35+-SQRT[1225+600])/2
X=(-35+-SQRT1825)/2
X=(-35+-42.72)/2
X=(-35+42.72)/2
X=7.72/2
X=3.86 ANSWER FOR THE BORDER.
PROOF
(30+2*3.86)(40+2*3.86)=1800
(30+7.72)(40+7.72)=1800
37.72*47.72=1800
1800=1800


Answer by renevencer22(21)

see
http://jrvencer3.esmartdesign.com/rdvmathematics.html
and look under algebra